# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v2

@@ 1. Introduction
To address these limitations, a promising strategy is to combine general-purpose but imprecise text prompts with dedicated interactions (e.g., clicks, drags, brushes, etc.) through graphical user interfaces (GUIs), as shown in Fig. 1-A. The synergy enables both flexible and fine-grained user control.   For example, consider Jason, a creative director designing a stylized magazine cover portrait. His intent is clear: “I need a stylized portrait for a magazine cover, with high fashion and artistic lighting”. Compared with the basic prompt-only workflow (Fig. 2-a), this task can be realized through a spectrum of interface designs (Fig. 2 b–f) that combine prompts and interactions in distinct ways. Concretely, some interfaces let him refine or expand the text prompt before generation (b), for instance, selecting “artistic lighting” for fine-tuning or switching to a “minimal black-and-white” style via UI controls. This functionality speeds up exploration but still assumes all intent can be expressed in language (Brade et al., 2023; Almeda et al., 2024). To go beyond language, Jason might manually sketch a portrait outline or brush key regions, merging these non-linguistic signals with the prompt for more precise spatial and referential control (c) (Lin et al., 2025; Kim et al., 2023b). When an initial render already exists, he may directly draw a necklace on the portrait, with the system incorporating both the prompt and the manipulated artifact content to enable precise, auditable edits grounded in the actual image (d) (Masson et al., 2024; Liu et al., 2025). After generating an initial image, some systems propose interactive controls (e), such as sliders for keyword (e.g., stylized” and “high fashion”) attention or color palettes for finer “artistic lighting”, allowing Jason to steer subsequent generations without rewriting text (Wang et al., 2024d; Gmeiner et al., 2025). Others analyze the generated artifact and propose structured options (f): Jason can choose to generate cover text, select the type, style, and length, and assemble a structured instruction for a magazine tagline (Zhang et al., 2023; Kim et al., 2023a).
@@
Building on this foundation, to examine if our IAI model can be generally applied to describe and differentiate existing GenAI systems’ interfaces, we revisit a curated corpus [^1] The annotation results: https://interaction-augmented-instruction.github.io/ of GenAI-powered interactive systems and map each system’s interaction workflow to one or more directed paradigm graphs (Fig. 1-C). From the corpus, we distill twelve frequent and composable atomic paradigms organized by interaction timing (pre- or post- GenAI invocation) and resource availability (prompt-only vs. artifact-grounded). These paradigm graphs based on the IAI model serve as high-level abstractions of interaction workflows, enabling tools to be clustered and compared.
@@ 3.2. Relations
The relation set is necessary because each retained relation reflects a non-reducible semantic flow, consistent with the principle of semantic meaningfulness. It is also sufficient because, after systematically enumerating and pruning all possible pairs, the remaining compositions span the full spectrum of prompt–interaction dynamics, aligning with discriminative value and agency preservation. We further demonstrate the descriptive, discriminative, and generative powers of our entity set and relation set through systematically review existing interface design and proposing new variations in Secs. 4 and 5.
@@ 3.3. Atomic Paradigm Graph
Fig. 3 depicts the full interaction-augmented instruction model as a directed graph. However, considering the task-specific needs and the complexity of integrating all entities and relations, not all of them must be adopted to represent an individual interface. Therefore, we introduce the concept of an atomic paradigm graph, defined as a minimal, self-contained subgraph of the model that captures a single, coherent interaction paradigm (Tab. 2). Here, atomic means that GenAI (G) plays only one role within the paradigm (avoiding overlapping or conflated functions) and that the paradigm necessarily involves interaction (I), as paradigms without interaction fall outside our scope. Each atomic paradigm graph is constructed by sequentially selecting entities and relations in the order they are enacted, thereby encoding the workflow through which a tool supports human–GenAI communication. Accordingly, a concrete tool can be represented by one or more atomic paradigm graphs, depending on how many distinct paradigms it supports. For example, an advanced image editing interface with brush-based region selection (Tab. 2-P4) would add H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug, and route the prompt through T $\rightarrow$ Aug $\rightarrow$ G, forming a richer paradigm (Masson et al., 2024; Liu et al., 2025). In another variant (Tab. 2-P8), if GenAI generates widgets to further tune generated artifacts, the graph will include G $\rightarrow$ I (GenAI initiates widget interactions) followed by H $\rightarrow$ I $\rightarrow$ A (human interacts with widgets to act on artifacts) (Vaithilingam et al., 2024). These small structural differences capture the paradigm characteristics.
@@ 4. Design Paradigms
To examine if our IAI model can be generally applied to describe and differentiate existing designs, we revisited and annotated 66 GenAI system interfaces that combine prompts and interactions in human-GenAI communications. We found that all workflows can be represented with atomic graphs derived from the model, demonstrating its descriptive capacity to capture recurring paradigms. Furthermore, we identified 12 recurring atomic interaction paradigms, which effectively capture the similarities and differences between interfaces. It verifies our IAI model’s power of differentiating interaction designs. We also hope that our identified paradigms can be a starting point for future interface design.
@@ 4.1. Revisit Corpora
Annotation was conducted by two authors using a structured, multi-stage protocol to ensure reliability and reproducibility. We began with a calibration pilot of 10 randomly sampled systems to test the initial codebook (grounded in IAI model design (Sec. 3) and definitions (Tab. 1), surface ambiguous cases, and refine operational rules. The two annotators then independently coded each system and cross-checked every annotation to reach a shared interpretation grounded in the codebook. All remaining uncertainties were reviewed and resolved with the involvement of a third author, and the resulting consensus annotations form the basis of our analysis. In addition to basic model definitions, our annotation followed these principles:
@@
To derive recurring paradigms, we canonicalized tool-specific labels into the standardized entities and relations and normalized away incidental UI differences that do not affect information flow. We then grouped the resulting atomic graphs by their topological structure to identify recurring interaction patterns. For each topological class, we generated a paradigm name, a canonical graph, a concise description, and a mapping to all systems exhibiting that structure. All tool metadata, annotation data, and grouping information are released to support replication222 The annotation results: https://interaction-augmented-instruction.github.io/.
@@
Interaction Before Calling GenAI Prompt-only (no artifact at hand) P1. Interactive Prompt Enhancement Human selects parts of drafted prompts for GenAI to refine or expand content. [Uncaptioned image] DreamSheets (Almeda et al., 2024), Promptify (Brade et al., 2023)
P2. Interactive Prompt Organization Human organize multiple text prompts into structured formats (e.g., tree) to fit the application. [Uncaptioned image] CoLadder (Yen et al., 2024), CoPrompt (Feng et al., 2024), PromptChainer (Wu et al., 2022), HoT (Nguyen et al., 2025), Intent Tagging (Gmeiner et al., 2025)
P3. Interaction as Part of Instruction Human’s interactions outside artifacts are included with prompts for GenAI to operate on artifacts. [Uncaptioned image] SketchFlex (Lin et al., 2025), InstructPipe (Zhou et al., 2025), Kim et al.  (Kim et al., 2023b)
Artifact-grounded P4. Referenced Artifact as Part of Instruction Human directly manipulates artifacts, which are combined with prompts as GenAI instructions. [Uncaptioned image] DirectGPT (Masson et al., 2024), Data Formulator (Wang et al., 2024a), ChatScratch (Chen et al., 2024), Qlarify (Fok et al., 2024), InternGPT (Liu et al., 2023b), MagicQuill (Liu et al., 2025), etc.
@@
Prompt-only (no artifact at hand) P5. AI-driven Prompt Suggestion GenAI suggests new or extended prompts from the human’s initial input for human selection. [Uncaptioned image] Spellburst (Angert et al., 2023), Intent Tagging (Gmeiner et al., 2025), PromptCharm (Wang et al., 2024d), Bursztyn et al.  (Bursztyn et al., 2021), PromptCrafter (Baek et al., 2023), AiCommentator (Andrews et al., 2024), Sensecape (Suh et al., 2023), ExploreLLM (Ma et al., 2024), etc.
@@
P7. Generative Prompt Control Widgets GenAI generates interactive widgets for fine-grained prompt control. [Uncaptioned image] PromptCharm (Wang et al., 2024d), DancingBoard (Chen et al., 2025b), Intent Tagging (Gmeiner et al., 2025)
@@
Artifact-grounded P9. Artifact to Structured Instruction GenAI uses artifacts as starting points to generate structured prompts. [Uncaptioned image] VISAR (Zhang et al., 2023), Metaphorian (Kim et al., 2023a), Intent Tagging (Gmeiner et al., 2025)
P10. Artifact to Multimodal Instruction GenAI parses and integrates artifacts to construct instructions. [Uncaptioned image] FigurA11y (Singh et al., 2024)
@@
P12. Interactive Artifact Refinement GenAI analyzes artifacts and initiates interactions based on prompts. [Uncaptioned image] PDFChatAnnotator (Tang et al., 2024)
@@ 4.2. Atomic Paradigms
Using the IAI model as an analytic lens, we identified 12 recurring atomic paradigms (Tab. 2) from our corpus. We organize these paradigms along two orthogonal dimensions based on our model: (1) interaction timing: whether interaction (I) occurs before or after invoking GenAI (G); and (2) user resources: whether the user begins interacting with GenAI (G) when they have no artifact (A) at hand (prompt-only) or are in the artifact-grounded situation. The timing dimension roughly tracks intent clarity and control locus: pre-invocation interactions are typical when users can specify constraints up front, whereas post-invocation interactions support exploratory or ambiguous goals via mixed-initiative refinement. The user resource dimension separates workflows where intent must be expressed solely in natural language from those where an existing artifact can be selected, annotated, or structured to ground the natural language instruction.
These dimensions yield four paradigm classes with characteristic graph topologies and consistent design applicability. Atomic paradigm graphs are designed to capture human–GenAI interaction logic and information flow, rather than to fully depict end-to-end system execution, and therefore do not necessarily terminate at the same node. For example, in pre-invocation paradigms, GenAI typically acts as a generator: users initiate interactions to construct an augmented instruction, which is then executed by GenAI, making the generated artifact a natural endpoint of the graph. In post-invocation paradigms, by contrast, GenAI often plays a clarifying or interpretive role: it inspects existing artifacts or prior generated results and initiates interactions to elicit user input, usually resulting in a refined instruction rather than a new artifact. Consistent with our annotation principles that each atomic paradigm graph contains only a single GenAI role, passing this refined instruction to a generator GenAI is represented as a separate atomic graph, and complete application workflows can be expressed through chains of such graphs. Paradigms that involve no interaction are outside our scope, as our focus is on characterizing interaction structures and information flow rather than all possible execution paths.
To better showcase the model’s descriptive and discriminative power, accordingly, the following subsections adopt a unified structure: we first describe the shared characteristics and typical cases of each paradigm class; we then analyze how similarities and differences in their paradigm graph structures map to concrete operations of each tool through representative comparisons (Fig. 4, Fig. 5, Fig. 6); and we conclude each subsection by articulating the key intra-class distinctions and the design implications they entail.
@@ 4.2.1. Pre-invocation, Prompt-only: Structuring and Enhancing Prompts Before GenAI Invocation
In the referenced artifact as part of instruction paradigm (P4), humans interact with artifacts to encapsulate either partial or entire content into augmented instruction (H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug; T $\rightarrow$ Aug). Systems such as DirectGPT (Masson et al., 2024) and MagicQuill (Liu et al., 2025) demonstrate this approach by passing dragged or brushed elements as precise constraints   (Fig. 4-P4 and Fig. 2-d). At first glance, DirectGPT could be mistaken for SketchFlex   (Fig. 4-P3 and Fig. 2-c) as they both create visual augmented instruction with interactions for GenAI to operate on artifacts. However, with our paradigm graphs, it is easy to notice that the key difference lies in how augmented instruction is composed: DirectGPT, as a representative of the referenced artifact as part of instruction paradigm, includes raw artifact segments, while SketchFlex, which belongs to the interaction as part of instruction paradigm category, encodes non-linguistic signals with interactions (H $\rightarrow$ I $\rightarrow$ Aug $\rightarrow$ G $\rightarrow$ A). This comparison highlights the key difference between the two paradigms: whether the interaction carries intent solely without embedding artifact.
@@ 4.2.3. Post-invocation, Prompt-only: Iterative Prompt Steering After GenAI Invocation
This class (Tab. 2 P5-P8) captures cases where users specify intent solely through prompts, but do so after an initial GenAI invocation. Unlike pre-invocation paradigms, where intent is clarified before execution, here the process is fundamentally iterative: the first prompt triggers model output, and subsequent refinement occurs in response to what GenAI returns. Because each atomic paradigm centers on a single GenAI role, these workflows often do not emphasize artifact creation, but rather focus on AI-assisted prompt steering and iterative negotiation of intent. This class arises when users’ initial instructions are vague, exploratory, or underspecified, and GenAI takes an active role in shaping subsequent prompts. Rather than users knowing exactly what they want, the system helps steer the process step by step.
@@
This class (Tab. 2, P9–P12) captures workflows in which an artifact—either produced by a prior GenAI call or provided directly by the user—serves as the anchor for subsequent interaction. Rather than relying solely on textual prompts, GenAI inspects the artifact, initiates interactions, and elicits user feedback to clarify or refine intent. Typical tasks include iteratively editing an image, interrogating a visualization, or correcting model-produced code.
The common structure is artifact-driven (A $\rightarrow$ G), with GenAI initiating interactions (G $\rightarrow$ I) that engage human responses (H $\rightarrow$ I). The four paradigms differ in how these interactions shape the construction of follow-up instructions. In artifact to structured instruction (P9), GenAI analyzes the artifact, generates a set of candidate operations or reformulations, and presents choices that users select to compose a structured textual prompt. For example, VISAR (Zhang et al., 2023) proposes expansion options for a selected paragraph; the user’s choices produce a structured prompt tree for subsequent text generation   (Fig. 4-P9 and Fig. 2-f). Building on this, artifact to multimodal instruction (P10) enriches the workflow by incorporating artifact itself. Here, GenAI extracts salient features from the artifact and asks users to tag or select them; the results are encoded into multimodal augmented instruction that combines artifact content with user intent as prompt. FigurA11y (Singh et al., 2024), for instance, extracts figure components and lets users link them to accessibility guidelines, producing a multimodal instruction (Fig. 6-P10). Moving beyond selection, artifact-driven prompt enhancement (P11) often positions the artifact as contextual grounding for semantic refinement. GenAI proposes contextual rules or mappings, which users confirm or modify in situ with interactions. LangAware (Chen et al., 2023) illustrates this by connecting low-level sensor signals to high-level contexts and enabling users to interactively combine contextual information with the original prompt before final execution (Fig. 6-P11). Finally, interactive artifact refinement (P12) emphasizes iterative analysis and correction. GenAI inspects the artifact in response to a prompt, surfaces candidate elements or annotations, and invites users to inspect, correct, or refine them. PDFChatAnnotator (Tang et al., 2024), for example, extracts information from PDFs and lets users guide annotation corrections interactively (Fig. 6-P12). Unlike prior paradigms, this focuses less on instruction construction and more on artifact-centered troubleshooting and refinement.
@@
Design situation (intent formulation) IAI-model rationale Paradigm
Fast, lightweight iterations on text Interaction only edits text (H→I→T), keeping instruction purely textual. P1
Need structured intent or task decomposition Interaction adds explicit structure to text (H→I→T→Aug). P2
Need non-linguistic or spatial expressivity Interaction supplements instruction (H→I→Aug, T→Aug) beyond text. P3
Reduce referential ambiguity via artifacts Artifact content grounds intent (I→A→Aug, T→Aug) for precise reference. P4
GenAI-assisted exploration after an initial output GenAI proposes refinements (G→I); users adjust via light edits (H→I→T). P5
@@
Need parameter-level control without editing text GenAI initiates parameter widgets (G→I) that augment prompt (H→I→Aug, T→Aug). P7
Need persistent, artifact-bound controls for iteration Generated widgets (G→I) attach to artifacts, enabling repeated edits (H→I→A). P8
Have artifacts but need help forming follow-up instructions GenAI analyzes the artifact to propose structured follow-ups (A→G→I, H→I→T→Aug). P9
Need multimodal follow-up instructions grounded in artifacts GenAI derives multimodal instruction candidates from artifacts (A→Aug). P10
Need to integrate artifact-derived suggestions into prompts GenAI proposes artifact-based actions (A→G→I) that supply prompts (H→I→Aug, T→Aug). P11

### v3

@@ 1. Introduction
To address these limitations, a promising strategy is to combine general-purpose but imprecise text prompts with dedicated interactions (e.g., clicks, drags, brushes, etc.) through graphical user interfaces (GUIs), as shown in Fig. 1-A. The synergy enables both flexible and fine-grained user control. For example, consider Jason, a creative director designing a stylized magazine cover portrait. His intent is clear: “I need a stylized portrait for a magazine cover, with high fashion and artistic lighting”. Compared with the basic prompt-only workflow (Fig. 2-A), this task can be realized through a spectrum of interface designs (Fig. 2 B–F) that combine prompts and interactions in distinct ways. Concretely, some interfaces let him refine or expand the text prompt before generation (Fig. 2-B), for instance, selecting “artistic lighting” for fine-tuning or switching to a “minimal black-and-white” style via UI controls. This functionality speeds up exploration but still assumes all intent can be expressed in language (Brade et al., 2023; Almeda et al., 2024). To go beyond language, Jason might manually sketch a portrait outline or brush key regions, merging these non-linguistic signals with the prompt for more precise spatial and referential control (Fig. 2-C) (Lin et al., 2025; Kim et al., 2023b). When an initial render already exists, he may directly draw a necklace on the portrait, with the system incorporating both the prompt and the manipulated artifact content to enable precise, auditable edits grounded in the actual image (Fig. 2-D) (Masson et al., 2024; Liu et al., 2025). After generating an initial image, some systems propose interactive controls (Fig. 2-E), such as sliders for keyword (e.g., stylized” and “high fashion”) attention or color palettes for finer “artistic lighting”, allowing Jason to steer subsequent generations without rewriting text (Wang et al., 2024d; Gmeiner et al., 2025). Others analyze the generated artifact and propose structured options (Fig. 2-F): Jason can choose to generate cover text, select the type, style, and length, and assemble a structured instruction for a magazine tagline (Zhang et al., 2023; Kim et al., 2023a).
@@
Building on this foundation, to examine if our IAI model can be generally applied to describe and differentiate existing GenAI systems’ interfaces, we revisit a curated corpus of GenAI-powered interactive systems and map each system’s interaction workflow to one or more directed paradigm graphs (Fig. 1-C). From the corpus, we distill twelve frequent and composable atomic paradigms organized by interaction timing (pre- or post-GenAI invocation) and resource availability (prompt-only vs. artifact-grounded) [^1] We provide an interactive browser for paradigms and annotated tools: https://interaction-augmented-instruction.github.io/. These paradigm graphs based on the IAI model serve as high-level abstractions of interaction workflows, enabling tools to be clustered and compared.
@@ 3.2. Relations
The relation set is necessary because each retained relation reflects a non-reducible semantic flow, consistent with the principle of semantic meaningfulness. It is also sufficient because, after systematically enumerating and pruning all possible pairs, the remaining compositions span the full spectrum of prompt–interaction dynamics, aligning with discriminative value and agency preservation. We further demonstrate the descriptive, discriminative, and generative powers of our entity set and relation set through systematically review existing interface design and proposing new variations in Sec. 4 and Sec. 5.
@@ 3.3. Atomic Paradigm Graph
Fig. 3 depicts the full interaction-augmented instruction model as a directed graph. However, considering task-specific needs and the complexity of integrating all entities and relations, not all of them must be adopted to represent an individual interface. Therefore, we introduce the concept of an atomic paradigm graph, defined as a minimal, self-contained subgraph of the model that captures a single, coherent interaction paradigm (Tab. 2). Here, atomic means that GenAI (G) plays only one role within the paradigm (avoiding overlapping or conflated functions) and that the paradigm necessarily involves Interaction (I), as paradigms without interactions fall outside our scope. Each atomic paradigm graph is constructed by sequentially selecting entities and relations in the order they are enacted, thereby encoding the workflow through which a tool supports human–GenAI communication. Accordingly, a concrete tool can be represented by one or more atomic paradigm graphs, depending on how many distinct paradigms it supports. For example, an advanced image editing interface with brush-based region selection (Tab. 2-P4) would add H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug, and route the prompt through T $\rightarrow$ Aug $\rightarrow$ G, forming a richer paradigm (Masson et al., 2024; Liu et al., 2025). In another variant (Tab. 2-P8), if GenAI generates widgets to further tune generated artifacts, the graph will include G $\rightarrow$ I (GenAI initiates widget interactions) followed by H $\rightarrow$ I $\rightarrow$ A (human interacts with widgets to act on artifacts) (Vaithilingam et al., 2024). These small structural differences capture paradigm characteristics.
@@ 4. Design Paradigms
To examine if our IAI model can be generally applied to describe and differentiate existing designs, we revisited and annotated 66 GenAI system interfaces that combine prompts and interactions in human-GenAI communications. We found that all workflows can be represented with atomic graphs derived from the model, demonstrating its descriptive capacity to capture recurring paradigms. Furthermore, we identified 12 recurring atomic interaction paradigms, which effectively capture similarities and differences between interfaces. It verifies our IAI model’s power of differentiating interaction designs. We also hope that our identified paradigms can be a starting point for future interface design.
@@ 4.1. Revisit Corpora
Annotation was conducted by two authors using a structured, multi-stage protocol to ensure reliability and reproducibility. We began with a calibration pilot of 10 randomly sampled systems to test the initial codebook (grounded in IAI model design (Sec. 3) and definitions (Tab. 1), surface ambiguous cases, and refine operational rules. The two annotators then independently coded half of the remaining systems and cross-checked each other’s annotation to reach a shared interpretation grounded in the codebook. All remaining uncertainties were reviewed and resolved with the involvement of a third author, and the resulting consensus annotations form the basis of our analysis. In addition to basic model definitions, our annotation followed these principles:
@@
To derive recurring paradigms, we canonicalized tool-specific labels into the standardized entities and relations and normalized away incidental UI differences that do not affect information flow. We then grouped the resulting atomic graphs by their topological structure to identify recurring interaction patterns. For each topological class, we generated a paradigm name, a canonical graph, a concise description, and a mapping to all systems exhibiting that structure. All tool metadata, annotation data, and grouping information are released to support replication in https://interaction-augmented-instruction.github.io/.
@@
Interaction Before Calling GenAI Prompt-only (no artifact at hand) P1. Interactive Prompt Enhancement A human selects parts of drafted text prompts for GenAI to refine or expand content. [Uncaptioned image] DreamSheets (Almeda et al., 2024), Promptify (Brade et al., 2023)
P2. Interactive Prompt Organization A human organizes multiple text prompts into structured formats (e.g., tree) to fit the application. [Uncaptioned image] CoLadder (Yen et al., 2024), CoPrompt (Feng et al., 2024), PromptChainer (Wu et al., 2022), HoT (Nguyen et al., 2025), IntentTagger (Gmeiner et al., 2025)
P3. Interaction as Part of Instruction A human’s interactions outside artifacts are included with prompts for GenAI to operate on artifacts. [Uncaptioned image] SketchFlex (Lin et al., 2025), InstructPipe (Zhou et al., 2025), Kim et al.  (Kim et al., 2023b)
Artifact-grounded P4. Referenced Artifact as Part of Instruction Human-selected partial or complete artifacts are combined with prompts as GenAI instructions. [Uncaptioned image] DirectGPT (Masson et al., 2024), Data Formulator (Wang et al., 2024a), ChatScratch (Chen et al., 2024), Qlarify (Fok et al., 2024), InternGPT (Liu et al., 2023b), MagicQuill (Liu et al., 2025), etc.
@@
Prompt-only (no artifact at hand) P5. AI-driven Prompt Suggestion GenAI suggests new or extended prompts from the human’s initial input for selection. [Uncaptioned image] Spellburst (Angert et al., 2023), IntentTagger (Gmeiner et al., 2025), PromptCharm (Wang et al., 2024d), Bursztyn et al.  (Bursztyn et al., 2021), PromptCrafter (Baek et al., 2023), AiCommentator (Andrews et al., 2024), Sensecape (Suh et al., 2023), ExploreLLM (Ma et al., 2024), etc.
@@
P7. Generative Prompt Control Widgets GenAI generates interactive widgets for fine-grained prompt control. [Uncaptioned image] PromptCharm (Wang et al., 2024d), DancingBoard (Chen et al., 2025b), IntentTagger (Gmeiner et al., 2025)
@@
Artifact-grounded P9. Artifact to Structured Instruction GenAI uses artifacts as starting points to generate structured instructions. [Uncaptioned image] VISAR (Zhang et al., 2023), Metaphorian (Kim et al., 2023a), IntentTagger (Gmeiner et al., 2025)
P10. Artifact to Multimodal Instruction GenAI parses and integrates artifacts to construct multimodal instructions. [Uncaptioned image] FigurA11y (Singh et al., 2024)
@@
P12. Interactive Artifact Refinement GenAI analyzes artifacts and initiates relevant interactions based on prompts. [Uncaptioned image] PDFChatAnnotator (Tang et al., 2024)
@@ 4.2. Atomic Paradigms
Using the IAI model as an analytic lens, we identified 12 recurring atomic paradigms (Tab. 2) from our corpus. We organize these paradigms along two orthogonal dimensions based on our model: (1) interaction timing: whether Interaction (I) occurs before or after invoking GenAI (G); and (2) user resources: whether the user begins interacting with GenAI (G) when they have no Artifact (A) at hand (prompt-only) or are in the artifact-grounded situation. The timing dimension roughly tracks intent clarity and control locus: pre-invocation interactions are typical when users can specify constraints up front, whereas post-invocation interactions support exploratory or ambiguous goals via mixed-initiative refinement. The user resource dimension separates workflows where intent must be expressed solely in natural language from those where an existing artifact can be selected, annotated, or structured to ground the natural language instruction. These dimensions yield four paradigm classes with characteristic graph topologies and consistent design applicability.
Atomic paradigm graphs are designed to capture human–GenAI interaction logic and information flow, rather than to fully depict end-to-end system execution, and therefore do not necessarily terminate at the same node. For example, in pre-invocation paradigms, GenAI typically acts as a generator: users initiate interactions to construct an augmented instruction, which is then executed by GenAI, making the generated artifact a natural endpoint of the graph. In post-invocation paradigms, by contrast, GenAI often plays a clarifying or interpretive role: it inspects existing artifacts or prior generated results and initiates interactions to elicit user input, usually resulting in a refined instruction rather than a new artifact. Consistent with our annotation principles that each atomic paradigm graph contains only a single GenAI role, passing this refined instruction to a generator GenAI is represented as a separate atomic graph, and complete application workflows can be expressed through chains of such graphs. Paradigms that involve no interaction are outside our scope, as our focus is on characterizing interaction structures and information flow rather than all possible execution paths.
To better showcase the model’s descriptive and discriminative power, accordingly, the following subsections adopt a unified structure: we first describe the shared characteristics and typical cases of each paradigm class; we then analyze how similarities and differences in their paradigm graph structures map to concrete operations of each tool through representative comparisons (Fig. 4, Fig. 5, and Fig. 6); and we conclude each subsection by articulating the key intra-class distinctions and the design implications they entail.
@@ 4.2.1. Pre-invocation, Prompt-only: Structuring and Enhancing Prompts Before GenAI Invocation
In the referenced artifact as part of instruction paradigm (P4), humans interact with artifacts to encapsulate either partial or entire content into augmented instruction (H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug; T $\rightarrow$ Aug). Systems such as DirectGPT (Masson et al., 2024) and MagicQuill (Liu et al., 2025) demonstrate this approach by passing dragged or brushed elements as precise constraints (Fig. 4-P4 and Fig. 2-D). At first glance, DirectGPT might look similar to SketchFlex (Fig. 4-P3 and Fig. 2-C), as they both allow humans to create visual augmented instruction with interactions for GenAI to operate on artifacts. However, with our paradigm graphs, it is easy to notice that the key difference lies in how augmented instruction is composed: DirectGPT, as a representative of the referenced artifact as part of instruction paradigm, includes raw artifact segments, while SketchFlex, which belongs to the interaction as part of instruction paradigm category, encodes non-linguistic signals with interactions (H $\rightarrow$ I $\rightarrow$ Aug $\rightarrow$ G $\rightarrow$ A). This comparison highlights the key difference between the two paradigms: whether the interaction carries intent solely without embedding artifact.
@@ 4.2.3. Post-invocation, Prompt-only: Iterative Prompt Steering After GenAI Invocation
This class (Tab. 2 P5-P8) captures cases where users specify intent through prompts and perform interactions to clarify intent after an initial GenAI invocation. Unlike pre-invocation paradigms, paradigms in this class often leverage GenAI for prompt steering and iterative negotiation of intent rather than directly instructing them for final artifact creation. This class arises when users’ initial instructions are vague, exploratory, or underspecified, and GenAI takes an active role in shaping subsequent prompts. Rather than users knowing exactly what they want, the system helps steer the process step by step.
@@
This class (Tab. 2, P9–P12) captures workflows in which an artifact serves as the anchor for subsequent interaction, regardless of whether the artifact is produced by a prior GenAI call or provided directly by the user. Rather than relying solely on textual prompts, GenAI inspects the artifact, initiates interactions, and elicits user feedback to clarify or refine intent. Typical tasks include iteratively editing an image, interrogating a visualization, or correcting model-produced code.
The common structure is artifact-driven (A $\rightarrow$ G), with GenAI initiating interactions (G $\rightarrow$ I) that engage human responses (H $\rightarrow$ I). The four paradigms differ in how these interactions shape the construction of follow-up instructions. In artifact to structured instruction (P9), GenAI analyzes the artifact, generates a set of candidate operations or reformulations, and presents choices that users select to compose a structured textual prompt. For example, VISAR (Zhang et al., 2023) proposes expansion options for a selected paragraph; the user’s choices produce a structured prompt tree for subsequent text generation (Fig. 4-P9 and Fig. 2-F). Building on this, artifact to multimodal instruction (P10) enriches the workflow by incorporating artifact itself. Here, GenAI extracts salient features from the artifact and asks users to tag or select them; the results are encoded into multimodal augmented instruction that combines artifact content with user intent as prompt. FigurA11y (Singh et al., 2024), for instance, extracts figure components and lets users link them to accessibility guidelines, producing a multimodal instruction (Fig. 6-P10). Moving beyond selection, artifact-driven prompt enhancement (P11) often positions the artifact as contextual grounding for semantic refinement. GenAI proposes contextual rules or mappings, which users confirm or modify in situ with interactions. LangAware (Chen et al., 2023) illustrates this by connecting low-level sensor signals to high-level contexts and enabling users to interactively combine contextual information with the original prompt before final execution (Fig. 6-P11). Finally, interactive artifact refinement (P12) emphasizes iterative analysis and improvement. GenAI inspects the artifact in response to a prompt, surfaces candidate elements or annotations, and invites users to inspect, correct, or refine them. PDFChatAnnotator (Tang et al., 2024), for example, extracts information from PDFs and lets users guide annotation corrections interactively (Fig. 6-P12). Unlike prior paradigms, this focuses less on instruction construction and more on artifact-centered troubleshooting and refinement.
@@
Design Situation Design Rationale Based on Our IAI Model Paradigm
Need fast, lightweight iterations on text prompts Interaction only edits text (H→I→T), keeping the instruction purely textual. P1
Need structured intent or task decomposition Interaction adds an explicit structure to text (H→I→T→Aug). P2
Need non-linguistic or spatial expressivity Interaction supplements the instruction (H→I→Aug, T→Aug) beyond text. P3
Need to reduce referential ambiguity via artifacts Artifact content grounds intent (I→A→Aug, T→Aug) for precise reference. P4
Need GenAI-assisted exploration after an initial output GenAI proposes refinements (G→I); users adjust via light edits (H→I→T). P5
@@
Need parameter-level control without editing text prompts GenAI initiates parameter widgets (G→I) that augment prompts (H→I→Aug, T→Aug). P7
Need persistent, artifact-bound controls for iteration Generated widgets (G→I) are attached to artifacts, enabling repeated edits (H→I→A). P8
Need to distill follow-up instructions from existing artifacts GenAI analyzes the artifact to propose structured follow-ups (A→G→I, H→I→T→Aug). P9
Need to form instructions with existing artifacts as references GenAI derives multimodal instruction candidates that incorporate artifacts (A→Aug). P10
Need artifact-derived suggestions to improve prompts GenAI proposes artifact-based actions (A→G→I) that supply prompts (H→I→Aug, T→Aug). P11